OpenStudy (anonymous):

Find the equation of the line passing through (4,3) and is a) parallel b) perpendicular to the line -3x+y-5=0

OpenStudy (mertsj):

-3x+y-5 = 0 has slope 3 The parallel line has the same slope so: y-3=3(x-4) The perpendicular has slope that is the negative reciprocal of 3 which is -1/3 So: y-3= -1/3(x-4)

OpenStudy (anonymous):

how do you that the line passing through (4,3) has a slope of 3 ? how do you find it ?

OpenStudy (anonymous):

How do we find the equation of the line passing through (4,3) ?

OpenStudy (mertsj):

The problem says that the line has to be parallel to -3x+y-5=0

OpenStudy (mertsj):

That line has slope = 3. Do you know that or not?

OpenStudy (anonymous):

yeah i know that -3x + y -5 =0 has a gradient of 3

OpenStudy (mertsj):

Well then a line parallel to it also has slope 3 and a line perpendicular to it has slope -1/3

OpenStudy (mertsj):

Do you understand now?

OpenStudy (anonymous):

Yes thanks a lot !

OpenStudy (mertsj):

yw